OpenStudy (anonymous):

Write the equation of the line with x-intercept of 3 and y-intercept of -5, in point-slope form.

OpenStudy (anonymous):

(y-y1) = m(x-x1)

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

ok, they give you both the x and y intercepts. this is two points on the line. x-intercept: (3,0) y-intercept: (0,-5) // this is just a point on the line. we are choosing the x-intercept y1 = 0 x1 = 3 so now we need to find slope which we can find using the slope formula. m = (y2-y1)/(x2-x1) m = (-5-0)/(0-3) m = -5/-3 m = 5/3 so we have (y-y1) = m(x-x1) (y-0) = (5/3)(x-3)

OpenStudy (ghazi):

 now simply see the two points and you are done using this equation \[(Y-Y _{1})=\frac{ (Y _{2}-Y _{1}) }{ (X _{2}-X _{1}) }*(X-X _{1})\] and X1,Y1 is (3,0) and X2,Y2 is (0,-5) , just plug in X2,X1 and Y2,Y1 , hope that helps @Megan_xoxo